A plumber can install a dishwasher in 50 minutes. If the plumber brings his apprentice to help, the job takes 40 minutes. How long would it take the apprentice working alone to install the dishwasher?
Found 2 solutions by oscargut, mangopeeler07:
Answer by oscargut(900)

In 40 minutes the plumber does 40/50 =0.8 of the work
so the apprentice does 0.2 of the work (20%)
then in 40 min does 20%
so 100% of the work is done in 200 min
Answer: 200 min

Answer by mangopeeler07(462)

A plumber can install a dishwasher in 50 minutes.
So plumber=1/50

because he installs 1/50 of a dishwasher in 1 minute

apprentice=x, because we don't know

plumber and apprentice=1/40

because together they install 1/40 of a dishwasher in 1 minute

1/50 + x/1=1/40

Common denominator
4/200 + 200x/200 = 5/200

4+200x=5

200x=1

x=1/200

So the apprentice can install 1/200 of a dishwasher in 1 minute
So in 200 minutes.

3 hours and 20 minutes